OpenStudy (anonymous):

Find the slope of the line that contains the following points. A(5, 6), B(10, 8)

OpenStudy (anonymous):

(5,6) (10,8) slope = (y2-y1)/(x2-x1) =(8-6)/(10-5) =2/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome :)